Open Question: The probability of selecting one person from the list who uses the Internet for less than 3 hours?
Open Question: Given a binomial probability experiment with six trials, in how many ways can we obtain two successes?
Open Question: Probability question ;D If two cards are drawn from a standard deck without replacement...?

Open Question: Probability question ;D If two cards are drawn from a standard deck without replacement...?

find the probability that at least two are red?

Help please! I got 1/6th for my answer o,O

Sorry, I could not read the content fromt this page.

View the original article here

Open Question: Probability problem with cards D:?

Open Question: Probability problem with cards D:?

I don't know much about cards so could someone help me?

If two cards are drawn from a standard deck of playing cards without replacement, find the probability that at least two are red?

I tried it out and I got 10/663 o,O isn't that way off? Could someone help me please!


View the original article here

Open Question: Probability help! From a collection of 5 blue and 4 red ink pens...?

Open Question: Probability help! From a collection of 5 blue and 4 red ink pens...?

three are selected at random. What is the probability that at least two are red?

I need some help! It'd be very much appreciated with learning how to do this :D!


View the original article here

Open Question: Help with genetics probability problem?

Open Question: Help with genetics probability problem?

Home >All Categories > Science & Mathematics >Mathematics >Open QuestionMusicalrain Musicalr... Member since:July 12, 2008Total points:2,753 (Level 4)Parents with the genotypes AaBbCc and AaBbCc are crossed, what is the probability that the offspring will have the genotype AaBbCc? I got 1/27 (3.7%) when I made a punnett square, but my prof wants us to calculate this like she did in class. I wasn't there for the class; what I can tell from the powerpoint is that she did this:
AaxAa=1/2 Aa
BbxBb=1/2 Bb
CcxCc=1/2 Cc
I'm not sure if she then calculated the probability like this:
(1/2)*(1/2)*(1/2)= 12.5%
Which method is correct? How would you calculate the probability by using fractions? I need to show my work, and there's no way that I can have a 27 square punnett square for each problem. Thanks!!Answer Question

Be the first to answer this question.


View the original article here

Open Question: Parents with the genotypes AaBbCc and AaBbCc are crossed, what is the probability that the offspring?

Open Question: Parents with the genotypes AaBbCc and AaBbCc are crossed, what is the probability that the offspring?

THe 1/2 X 1/2 X 1/2 = 1/8 = 0.125 is correct

You are looking at each independently-assorting trait independently. Using the multiplication rule of probability, probability of 3 independent events occurring simultaneously is the product of their individual probabilities.

In your Punnet Square, you would have an 8X8 square with 64 boxes. 8 of those should have had AaBbCc in them, giving you a 1/8 probability.


View the original article here

Open Question: What is the probability of rolling a sum of 11 with two dice?

Open Question: What is the probability of rolling a sum of 11 with two dice?

While I haven't taken a statistics class, it seems like the answer would be 2 out of 36 (or 1 out of 18).

There are 36 possible outcomes when rolling two dice (for each of the 6 possibilities on die A, there are 6 possibilities on die B). Only two of these outcomes (6,5) and (5,6) add up to 11.


View the original article here

Open Question: Do you have to be proficient in probability and statistics to be a quantum mechanical physicist?
Open Question: Probability In Math Question?

Open Question: Probability In Math Question?

Home >All Categories > Science & Mathematics >Mathematics >Open QuestionScott Scott Member since:February 27, 2011Total points:52 (Level 1)The incoming freshman class at Winston College shows that 30% come from single- parent homes and 62% come from two- parent homes. The Financial aid office reports that 65% of students from single-parent homes recieve finanical aid and 42% of those from two-parent homes receive financial aid.
Find the Probability that an incoming freshman who recieves financial sid is from the two-parent home????
Can someone help me figure this question out!!! Thanks so much..Answer Question

Be the first to answer this question.


View the original article here

Open Question: Help with probability problem please?

Open Question: Help with probability problem please?

Cars are being produced by two factories, but factory I produces twice as many cars as factory II in a given time. Factory I is known to produce 1% defectives and factory II produces 3% defectives. A car is examined and found to be defective. What is the probability the car came from factory I if it is defective?

Please show all work I have no idea where to begin. Thank you.


View the original article here

Open Question: Probability: investments (possibly a Venn Diagram)?

Open Question: Probability: investments (possibly a Venn Diagram)?

In a survey of 200 employees of a company
regarding their 401(k) investments, the following data
were obtained:
141 had investments in stock funds.
91 had investments in bond funds.
60 had investments in money market funds.
47 had investments in stock funds and bond funds.
36 had investments in stock funds and money market
funds.
36 had investments in bond funds and money market
funds.
22 had investments in stock funds, bond funds, and
money market funds.

What is the probability that an employee of the company
chosen at random
a. Had investments in exactly two kinds of investment
funds?
b. Had investments in exactly one kind of investment
fund?
c. Had no investment in any of the three types of funds?


View the original article here

Open Question: probability math question?

Open Question: probability math question?

10W
5R

1/10*1/9
+
1/5*1/4

I believe, i think thats wrong though

Sorry, I could not read the content fromt this page.

View the original article here